# MolTC: Towards Molecular Relational Modeling In Language Models

### v5

@@ Abstract
Molecular Relational Learning (MRL), aiming to understand interactions between molecular pairs, plays a pivotal role in advancing biochemical research. Recently, the adoption of large language models (LLMs), known for their vast knowledge repositories and advanced logical inference capabilities, has emerged as a promising way for efficient and effective MRL. Despite their potential, these methods predominantly rely on textual data, thus not fully harnessing the wealth of structural information inherent in molecular graphs. Moreover, the absence of a unified framework exacerbates the issue of insufficient data exploitation, as it hinders the sharing of interaction mechanism learned across various datasets. To address these challenges, this work proposes a novel LLM-based multi-modal framework for Molecular inTeraction modeling following Chain-of-Thought (CoT) theory, termed MolTC, which effectively integrate graphical information of two molecules in pair. For achieving a unified training paradigm, MolTC innovatively develops a Dynamic Parameter-sharing Strategy for cross-dataset information exchange. Moreover, to train this integrated framework efficiently, we introduce a Multi-hierarchical CoT theory to refine its training paradigm, and conduct a comprehensive Molecular Interactive Instructions dataset for the development of biochemical LLMs involving MRL. Our experiments, conducted across twelve datasets involving over 4,000,000 molecular pairs, exhibit the superiority of our method over current GNN and LLM-based baselines. Code is available at https://github.com/MangoKiller/MolTC.
@@ 1 Introduction
To overcome these limitations, in this work, we propose MolTC, a unified multi-modal framework for Molecular inTeraction modeling following the Chain-of-thought theory Wei et al. (2022). As depicted in Figure 1 (c), MolTC employs the Graph Neural Networks (GNNs) Kipf and Welling (2017), known for their proficiency in graph modeling, to explicitly gather graphical information of molecular pairs, and integrates them into the input space of LLMs by two meticulously crafted projectors. In response to empirical findings that LLMs may confuse two input molecules in pair, MolTC incorporates the molecules’ SMILES information to reinforce the concept of molecular order. More importantly, to achieve a unified learning paradigm, MolTC develops a Dynamic Parameter-sharing strategy for bolstering cross-dataset information exchange, which can boost the efficiency and effectiveness simultaneously.
Based on these, a two-pronged approach is developed to train this integrated framework efficiently:
@@ 2.1 Framework of MolTC
Here we introduce four key components of MolTC’s framework: Graph Encoder, Representation Projector, SIMLES Injector, and the backbone LLM. The specific instantiation details of each module can be found in the experimental section and the appendix.
@@
Backbone LLM. MolTC leverages Galactica, a decoder-only transformer built on the OPT framework, as its backbone LLM. Pretrained on an extensive collection of scientific literature, Galactica demonstrates exceptional proficiency in biochemistry knowledge. This expertise, particularly in parsing molecular sequences such as SMILES and SELFIES strings, enables Galactica to adeptly capture the properties crucial for molecular interactions. Specifically, the goal of MolTC is to harness Galactica’s advanced inferential skills to interpret the contextual interactions between two molecular sets of token collections, $\{\mathbf{M}_{a},\mathbf{S}_{a}\}$ and $\{\mathbf{M}_{b},\mathbf{S}_{b}\}$. More formally, we denote an integrated prompt sequence as follows:
@@ 3.1 Experimental Setting
Baselines. For a comprehensive evaluation, we conduct various baseline methods encompassing distinct categories such as methods based on: GNNs, DL models other than GNN, and LLMs. Specifically, For DDI task, we employ GoGNN Wang et al. (2020), MHCADDI Deac et al. (2019), DeepDDI Ryu et al. (2018), SSI-DDI, CGIB Lee et al. (2023a), CMRL Lee et al. (2023b), MDF-SA-DDI Lin et al. (2022), DSN-DDI Li et al. (2023c) as the backbone. For SSI and CSI tasks, we utilize D-MPNN Vermeire and Green (2021), SolvBert Yu et al. (2023), SMD Meng et al. (2023), CGIB Lee et al. (2023a), CIGIN Pathak et al. (2020), GEM Fang et al. (2022), GOVER Rong et al. (2020), Uni-Mol Zhou et al. (2023) as the backbone. Furthermore, all downstream tasks adopt LLM-based methods, such as Galactica Taylor et al. (2022), Chem T5 Christofidellis et al. (2023), MolT5 Edwards et al. (2022) and MolCA Liu et al. (2023) as the backbone.
@@ Appendix A Related Work
Since exhaustive experimental validation of the molecule interactions is notoriously time-consuming and costly Lee et al. (2023a), more recently, adopting LLM has emerged as a promising alternative for efficient and effective molecular relational learning, which are known for their vast knowledge repositories and advanced logical inference capabilities. Specifically,
  • Park et al. (2022); Jha et al. (2022a, b) focus on employ LLM to optimize protein-protein interactions (PPI) tasks. In this context, proteins are represented as residue contact graphs, also known as amino acid graphs, where each node is a residue Jha et al. (2022b). Notably, Jha et al. (2022b) leverages the superior encoding capabilities of the biochemical LLMs, where the input to the LLM is the protein sequence, and the output is a feature vector for each amino acid in the sequence. This output is then used as node features in the residue contact graph to enhance the prediction of PPI tasks.
  • Sagawa and Kojima (2023); Chen et al. (2023); Livne et al. (2023); Shi et al. (2023) focus using LLMs to optimize chemical reactions. Specifically, Shi et al. (2023) selects in-context reaction examples with varying confidence scores closest to the target reaction query, encouraging large models to understand the relationships between these reactions. Sagawa and Kojima (2023) focuses on optimizing low-sample organic chemical applications by pretraining them with extensive compound libraries and fine-tuning with smaller in-house datasets for specific tasks. Livne et al. (2023) introduces a new foundational model, nach0, capable of solving various chemical and biological tasks, including molecular synthesis.
  • Li et al. (2023b); Pei et al. (2023) focus on using LLMs to optimize tasks related to drug molecules. Specifically, Pei et al. (2023) enriches cross-modal integration in biology with chemical knowledge and natural language associations, achieving significant results in multiple drug-target interaction prediction tasks. Meanwhile, Li et al. (2023b) concentrates on few-shot drug pair synergy prediction.
Furthermore, it is noteworthy that due to the significance of molecular tasks such as DDI for biochemical LLMs Fang et al. (2023d), the Mol-instructions dataset has been introduced more recently Fang et al. (2023c). This dataset shares many similarities with our work, and can be considered as complementary to each other.
@@ B.1 Datasets
ZhangDDI. Zhang et al. (2017) it contains 548 drugs and 48,548 pairwise interaction data and multiple types of similarity information about these drug pairs.
ChChMiner. Zitnik et al. (2018) it contains 1,322 drugs and 48,514 labeled DDIs, obtained through drug labels andscientific publications.
DeepDDI. Ryu et al. (2018) contains 192,284 labeled DDIs and their detailed side-effect information, which is extracted from Drugbank.
TWOSIDES. Tatonetti et al. (2012) it collected 555 drugs and their 3,576,513 pairwise interactions involving 1318 interaction types from TWOSIDES.
Chromophore. Joung et al. (2020) contains 20,236 combinations of 7,016 chromophores and 365 solvents which are given in the SMILES string format. All optical properties are based on scientific publications and unreliable experimental results are excluded after examination of absorption and emission spectra. In this dataset, we measure our model performance on predicting maximum absorption wavelength (Absorption), maximum emission wavelength (Emission) and excited state lifetime (Lifetime) properties which are important parameters for the design of chromophores for specific applications. We delete the NaN values to create each dataset which is not reported in the original scientific publications. Moreover, for Lifetime data, we use log normalized target value since the target value of the dataset is highly skewed inducing training instability.
MNSol. Marenich et al. (2020) contains 3,037 experimental free energies of solvation or transfer energies of 790 unique solutes and 92 solvents.In this work, we consider 2,275 combinations of 372 unique solutes and 86 solvents following previous work.
FreeSolv. Mobley and Guthrie (2014) provides 643 experimental and calculated hydration free energy of small molecules in water. In this work, we consider 560 experimental results following previous work.
CompSol. Moine et al. (2017) dataset is proposed to show how solvation energies are influenced by hydrogen-bonding association effects. We consider 3,548 combinations of 442 unique solutes and 259 solvents in the dataset following previous work.
Abraham. Grubbs et al. (2010) dataset is a collection of data published by the Abraham research group at College London. We consider 6,091 combinations of 1,038 unique solutes and 122 solvents following previous work.
CombiSolv. Vermeire and Green (2021) contains all the data of MNSol, FreeSolv, CompSol, and Abraham, resulting in 10,145 combinations of 1,368 solutes and 291 solvents.
CombiSolv-QM. Vermeire and Green (2021) is generated with 1 million combinations of 284 commonly used solvents and 11,029 solutes. Those 1 million data points are randomly selected from all possible solvent–solute combinations. Solvents and solutes with elements $H,B,C,N,O,F,P,S,Cl,Br$ and $I$ are included with a solute molar mass ranging from 2.02 g/mol to 1776.89 g/mol.
@@ B.2 Baselines
GoGNN. Wang et al. (2020) It extracts features from structured entity graphs and entity interaction graphs in a hierarchical manner. We also propose a dual attention mechanism that enables the model to preserve the importance of neighbors in both levels of the graph.
MHCADDI. Deac et al. (2019) A gated information transfer neural network is used to control the extraction of substructures and then interact based on an attention mechanism.
DeepDDI. Ryu et al. (2018) First, the structural similarity profile is calculated between the two input drugs and other drugs, and then prediction is completed based on the deep neural network.
SSI-DDI. Nyamabo et al. (2021) it use a 4-layer GAT network to extract substructures at different levels, and finally complete the final prediction based on the co-attention mechanism
CGIB. Lee et al. (2023a) Based on the graph conditional information bottleneck theory, conditional subgraphs are extracted to complete the interaction between molecules.
CMRL. Lee et al. (2023b) it detects the core substructure that is causally related to chemical reactions. we introduce a novel conditional intervention framework whose intervention is conditioned on the paired molecule. With the conditional intervention framework.
MDF-SA-DDI. Lin et al. (2022) it predicts interaction (DDI) events based on multi-source drug fusion, multi-source feature fusion and transformer self-attention mechanism.
DSN-DDI. Li et al. (2023c) it employs local and global representation learning modules iteratively and learns drug substructures from the single drug ‘intra-view’) and the drug pair (‘inter-view’) simultaneously.
@@
D-MPNN Vermeire and Green (2021) it employes a transfer learning approach to predict solvation free energies, integrating quantum calculation fundamentals with the heightened accuracy of experimental measurements through two new databases, CombiSolv-QM and CombiSolv-Exp.
SolvBert. Yu et al. (2023) it interprets solute and solvent interactions through their combined SMILES representation. Pre-trained using unsupervised learning with a substantial computational solvation free energy database, SolvBERT is adaptable to predict experimental solvation free energy or solubility by fine-tuning on specific databases.
SMD. Meng et al. (2023) utilizes the quantum charge density of a solute and a continuum representation of the solvent. It breaks down solvation free energy into two components: bulk electrostatic contribution, treated through a self-consistent reaction field using IEF-PCM, and a cavity-dispersion-solvent-structure term, accounting for short-range interactions in the solvation shell based on atomic surface areas with geometry-dependent constants.
CIGIN. Pathak et al. (2020) is a method based on graph neural networks. The proposed model adopts an end-to-end framework consisting of three essential phases: message passing, interaction, and prediction. In the final phase, these stages are leveraged to predict solvation free energies.
GEM. Fang et al. (2022) exhibits a uniquely designed geometry-based graph neural network architecture, complemented by several dedicated self-supervised learning strategies at the geometry level. That aims to acquire comprehensive molecular geometry knowledge for accurate prediction of molecular properties.
GOVER. Rong et al. (2020) captures rich structural information from extensive unlabeled molecular data through self-supervised tasks, employing a flexible Transformer-style architecture integrated with Message Passing Networks. This allows GROVER to be trained efficiently on large-scale datasets without supervision, addressing data scarcity and bias challenges.
Uni-Mol.Zhou et al. (2023) incorporates two pre-trained models featuring the SE(3) Transformer architecture: a molecular model pre-trained on 209 million molecular conformations and a pocket model pre-trained on 3 million candidate protein pocket data. Additionally, Uni-Mol integrates various fine-tuning strategies to effectively apply these pre-trained models across diverse downstream tasks.

### v6

@@ Abstract
Molecular Relational Learning (MRL), aiming to understand interactions between molecular pairs, plays a pivotal role in advancing biochemical research. Recently, the adoption of large language models (LLMs), known for their vast knowledge repositories and advanced logical inference capabilities, has emerged as a promising way for efficient and effective MRL. Despite their potential, these methods predominantly rely on textual data, thus not fully harnessing the wealth of structural information inherent in molecular graphs. Moreover, the absence of a unified framework exacerbates the issue of insufficient data exploitation, as it hinders the sharing of interaction mechanisms learned across various datasets. To address these challenges, this work proposes a novel LLM-based multi-modal framework for Molecular inTeraction modeling following Chain-of-Thought (CoT) theory, termed MolTC, which effectively integrate graphical information of two molecules in pair. To train MolTC efficiently, we introduce a Multi-hierarchical CoT theory to refine its training paradigm, and conduct a comprehensive Molecular Interactive Instructions dataset for the development of biochemical LLMs involving MRL. Our experiments, conducted across twelve datasets involving over 4,000,000 molecular pairs, exhibit the superiority of our method over current GNN- and LLM-based baselines. Our code is available at https://github.com/MangoKiller/MolTC.
@@ 1 Introduction
To overcome these limitations, in this work, we propose MolTC, a unified multi-modal framework for Molecular inTeraction modeling following the Chain-of-thought theory Wei et al. (2022). As depicted in Figure 1 (c), MolTC employs the Graph Neural Networks (GNNs) Kipf and Welling (2017), known for their proficiency in graph modeling, to explicitly gather graphical information of molecular pairs, and integrates them into the input space of LLMs by two meticulously crafted projectors. In response to empirical findings that LLMs may confuse two input molecules in pair, MolTC incorporates the molecules’ SMILES information to reinforce the concept of molecular order.
More importantly, a two-pronged approach is developed to train MolTC efficiently:
@@ 2.1 Framework of MolTC
Here we introduce four key components of MolTC’s framework: Graph Encoder, Representation Projector, SIMLES Injector, and the backbone LLM. The specific instantiation details can be found in the experimental section and the appendix.
@@
Backbone LLM. MolTC leverages Galactica, a decoder-only transformer built on the OPT framework, as its base LLM. Pretrained on an extensive collection of scientific literature, Galactica demonstrates exceptional proficiency in biochemistry knowledge. This expertise, particularly in parsing molecular sequences such as SMILES and SELFIES strings, enables Galactica to adeptly capture the properties crucial for molecular interactions. Specifically, the goal of MolTC is to harness Galactica’s advanced inferential skills to interpret the contextual interactions between two molecular sets of token collections, $\{\mathbf{M}_{a},\mathbf{S}_{a}\}$ and $\{\mathbf{M}_{b},\mathbf{S}_{b}\}$. More formally, we denote an integrated prompt sequence as follows:
@@ 3.1 Experimental Setting
Baselines. For a comprehensive evaluation, we conduct various baseline methods encompassing distinct categories such as methods based on: GNNs, DL models other than GNN, and LLMs. Specifically, For DDI task, we employ GoGNN Wang et al. (2020), MHCADDI Deac et al. (2019), DeepDDI Ryu et al. (2018), SSI-DDI, CGIB Lee et al. (2023a), CMRL Lee et al. (2023b), MDF-SA-DDI Lin et al. (2022), DSN-DDI Li et al. (2023c) as the backbone. For SSI and CSI tasks, we utilize D-MPNN Vermeire and Green (2021), SolvBert Yu et al. (2023a), SMD Meng et al. (2023), CGIB Lee et al. (2023a), CIGIN Pathak et al. (2020), GEM Fang et al. (2022), GOVER Rong et al. (2020), Uni-Mol Zhou et al. (2023) as the backbone. Furthermore, all downstream tasks adopt LLM-based methods, such as Galactica Taylor et al. (2022), Chem T5 Christofidellis et al. (2023), MolT5 Edwards et al. (2022) and MolCA Liu et al. (2023a) as the backbone.
@@ 3.4 Ablation Study (RQ3)
4 Discussion
Potential Data Leakage Risk. LLM might have memorized the prediction target during pretraining using a vast amount of text corpus. Hence, we acknowledge that LLM-based methods for biochemistry tasks such as MolT5 Edwards et al. (2022), MolCA Liu et al. (2023a) and our MolTC may face potential data leakage risks due to the memory during the pre-training process of their base LLMs. However, we claim that ablation experiments can mitigate the influence of data leakage, demonstrating the effectiveness of methods despite using potentially contaminated LLMs. For instance, in our ablation experiments, the backbone LLMs are fed with SMILES strings and textual descriptions of molecules solely, and their performance is much worse than that of original methods. This stark contrast leads to two analytical observations:
  • Firstly, note that these text inputs, highly recognizable and scattered throughout the LLM’s training data, could provide all the necessary information to identify a molecule. Hence, if there is significant data leakage, the LLM could easily recall from memory and make reasonably accurate predictions based on these inputs. Based on this, the poor performance of these ablation experiments can, to some extent, counter the concern of data leakage.
  • Secondly, since these ablation experiments and the original experiments utilized the exact same backbone LLM, the significant differences in their outcomes indicate that the original method’s design is highly effective, regardless of whether the backbone LLM is subject to data leakage.
In summary, while these ablation experiments can not completely eliminate the impact of data leakage, they do mitigate it. Furthermore, it is worth noting that while the risk of data leakage exists, it does not deter researchers from continuously exploring and optimizing LLM-based frameworks for biochemical tasks. In the future, by constructing new biochemical datasets through biochemical experiments, which do not exist in the LLM’s training data, researchers can eliminate the data leakage issue in these frameworks thoroughly.
@@ 5 Conclusion
Acknowledgement
This research is supported by the National Science and Technology Major Project (2023ZD0121102), National Natural Science Foundation of China (92270114).
@@ Appendix A Related Work
Since exhaustive experimental validation of the molecule interactions is notoriously time-consuming and costly Lee et al. (2023a), more recently, adopting LLM has emerged as a promising alternative for efficient and effective molecular relational learning, which are known for their vast knowledge repositories and advanced logical inference capabilities. Compared to the prediction of single-molecule properties Liu et al. (2024b); Li et al. (2024); Liu et al. (2023b); Fang et al. (2023d); Guo et al. (2023a), the field of molecular interaction prediction is still in its early stages. For instance,
  • Protein-protein interactions (PPI). In this context, proteins are represented as residue contact graphs, also known as amino acid graphs, where each node is a residue Park et al. (2022); Jha et al. (2022a, b, b). Notably, Jha et al. (2022b) leverages the superior encoding capabilities of the biochemical LLMs, where the input to the LLM is the protein sequence, and the output is a feature vector for each amino acid in the sequence. This output is then used as node features in the residue contact graph to enhance the prediction of PPI tasks.
  • DDI. In this context, Pei et al. (2023) focuses on enriching cross-modal integration in biology with chemical knowledge and natural language associations, achieving significant results in multiple drug-target interaction prediction tasks. Recently, few-shot drug pair synergy prediction is gradually gaining attention and entering the spotlightLi et al. (2023b).
  • Chemical reactions. For understanding chemical reactions Chen et al. (2023); Guo et al. (2023b), Shi et al. (2023) selects in-context reaction examples with varying confidence scores closest to the target reaction query, encouraging large models to understand the relationships between these reactions. Sagawa and Kojima (2023) focuses on optimizing low-sample organic chemical applications by pretraining them with extensive compound libraries and fine-tuning with smaller in-house datasets for specific tasks. Livne et al. (2023) introduces a new foundational model, nach0, capable of solving various chemical and biological tasks, including molecular synthesis.
More recently, ReactXT Liu et al. (2024a) also focuses on modeling molecular interactions using chemical reactions. However, it is unable to leverage CoT for more robust information extraction. Meanwhile, the Mol-instructions dataset has been introduced Fang et al. (2023c) for effectively training biochemistry LLMs, that can be considered as complementary to our MoT-instruction.
@@ B.1 Datasets
ZhangDDI Zhang et al. (2017). It contains 548 drugs and 48,548 pairwise interaction data and multiple types of similarity information about these drug pairs.
ChChMiner Zitnik et al. (2018). It contains 1,322 drugs and 48,514 labeled DDIs, obtained through drug labels andscientific publications.
DeepDDI Ryu et al. (2018). It contains 192,284 labeled DDIs and their detailed side-effect information, which is extracted from Drugbank.
TWOSIDES Tatonetti et al. (2012). It collects 555 drugs and their 3,576,513 pairwise interactions involving 1318 interaction types from TWOSIDES.
Chromophore Joung et al. (2020). It contains 20,236 combinations of 7,016 chromophores and 365 solvents which are given in the SMILES string format. All optical properties are based on scientific publications and unreliable experimental results are excluded after examination of absorption and emission spectra. In this dataset, we measure our model performance on predicting maximum absorption wavelength (Absorption), maximum emission wavelength (Emission) and excited state lifetime (Lifetime) properties which are important parameters for the design of chromophores for specific applications. We delete the NaN values to create each dataset which is not reported in the original scientific publications. Moreover, for Lifetime data, we use log normalized target value since the target value of the dataset is highly skewed inducing training instability.
MNSol Marenich et al. (2020). It contains 3,037 experimental free energies of solvation or transfer energies of 790 unique solutes and 92 solvents.In this work, we consider 2,275 combinations of 372 unique solutes and 86 solvents following previous work.
FreeSolv Mobley and Guthrie (2014). It provides 643 experimental and calculated hydration free energy of small molecules in water. In this work, we consider 560 experimental results following previous work.
CompSol Moine et al. (2017). This dataset is proposed to show how solvation energies are influenced by hydrogen-bonding association effects. We consider 3,548 combinations of 442 unique solutes and 259 solvents in the dataset following previous work.
Abraham Grubbs et al. (2010). This dataset is a collection of data published by the Abraham research group at College London. We consider 6,091 combinations of 1,038 unique solutes and 122 solvents following previous work.
CombiSolv Vermeire and Green (2021). It contains all the data of MNSol, FreeSolv, CompSol, and Abraham, resulting in 10,145 combinations of 1,368 solutes and 291 solvents.
CombiSolv-QM Vermeire and Green (2021). It is generated with 1 million combinations of 284 commonly used solvents and 11,029 solutes. Those 1 million data points are randomly selected from all possible solvent–solute combinations. Solvents and solutes with elements $H,B,C,N,O,F,P,S,Cl,Br$ and $I$ are included with a solute molar mass ranging from 2.02 g/mol to 1776.89 g/mol.
@@ B.2 Baselines
GoGNN Wang et al. (2020). It extracts features from structured entity graphs and entity interaction graphs in a hierarchical manner. We also propose a dual attention mechanism that enables the model to preserve the importance of neighbors in both levels of the graph.
MHCADDI Deac et al. (2019). A gated information transfer neural network is used to control the extraction of substructures and then interact based on an attention mechanism.
DeepDDI Ryu et al. (2018). First, the structural similarity profile is calculated between the two input drugs and other drugs, and then prediction is completed based on the deep neural network.
SSI-DDI Nyamabo et al. (2021). It uses a 4-layer GAT network to extract substructures at different levels, and completes the final prediction based on the co-attention mechanism.
CGIB Lee et al. (2023a). Based on the graph conditional information bottleneck theory, conditional subgraphs are extracted to complete the interaction between molecules.
CMRL Lee et al. (2023b). It detects the core substructure that is causally related to chemical reactions by introducing a novel conditional intervention framework whose intervention is conditioned on the paired molecule.
MDF-SA-DDI Lin et al. (2022). It predicts interaction based on multi-source drug fusion, multi-source feature fusion and transformer self-attention mechanism.
DSN-DDI Li et al. (2023c). It employs local and global representation learning modules iteratively and learns drug substructures from the single drug ‘intra-view’) and the drug pair (‘inter-view’) simultaneously.
@@
D-MPNN Vermeire and Green (2021). It employs a transfer learning approach to predict solvation free energies, integrating quantum calculation fundamentals with the heightened accuracy of experimental measurements through two new databases, CombiSolv-QM and CombiSolv-Exp.
SolvBert Yu et al. (2023a). It interprets solute and solvent interactions through their combined SMILES representation. After pre-training by unsupervised learning with a substantial computational solvation free energy database, SolvBERT is adaptable to predict experimental solvation free energy or solubility by fine-tuning on specific databases.
SMD Meng et al. (2023). It utilizes the quantum charge density of a solute and a continuum representation of the solvent. It breaks down solvation free energy into two components: bulk electrostatic contribution, treated through a self-consistent reaction field using IEF-PCM, and a cavity-dispersion-solvent-structure term, accounting for short-range interactions in the solvation shell based on atomic surface areas with geometry-dependent constants.
CIGIN Pathak et al. (2020). It adopts an end-to-end framework consisting of three essential phases: message passing, interaction, and prediction. In the final phase, these stages are leveraged to predict solvation free energies.
GEM Fang et al. (2022). It exhibits a uniquely designed geometry-based graph neural network architecture, complemented by several dedicated self-supervised learning strategies at the geometry level. That aims to acquire comprehensive molecular geometry knowledge for accurate prediction of molecular properties.
GOVER Rong et al. (2020). It captures rich structural information from extensive unlabeled molecular data through self-supervised tasks, employing a flexible Transformer-style architecture integrated with Message Passing Networks. This allows GROVER to be trained efficiently on large-scale datasets without supervision, addressing data scarcity and bias challenges.
Uni-Mol Zhou et al. (2023). It incorporates two pre-trained models featuring the SE(3) Transformer architecture: a molecular model pre-trained on 209 million molecular conformations and a pocket model pre-trained on 3 million candidate protein pocket data. Additionally, Uni-Mol integrates various fine-tuning strategies to effectively apply these pre-trained models across diverse downstream tasks.